# Interpolating neural network: A lightweight yet precise architecture for data training, equation solving, and parameter calibration

### v4

@@ 3 Scientific and Engineering applications: Metal Additive Manufacturing
This section illustrates how Engineering Software 2.0 enabled with INNs can be applied to advance various facets of metal additive manufacturing (AM). Note that the INN with CP decomposition will be referred to as INN in this section for brevity. Laser powder bed fusion (L-PBF) is a type of metal additive manufacturing in which metal powders are layered and melted by a laser beam to create objects that match the provided CAD designs [47]. The added flexibility of L-PBF comes with a huge design space and spatially varying microstructure after manufacturing. Therefore, considerable research is devoted to employing in situ monitoring data for real-time control of L-PBF processes to minimize the variability and optimize the target properties [48]. Computational modeling of L-PBF spans the simulation of the manufacturing process with a high dimensional design space (solving), the identification of model parameters from sparse experimental data (calibrating), and the online monitoring and control of the L-PBF (training).
@@ 3.1 Real-time Online Control of Laser Powder Bed Fusion Additive Manufacturing
The first application is fully data-driven, where the data are generated from numerical methods and the INN trainer is used to develop an online control system for L-PBF AM. The goal is to maintain a homogeneous temperature of the melt pool (i.e., a molten region near the laser spot) across each layer using real-time sensor data, under the hypothesis that a homogeneous melt pool temperature across each layer of material deposition will reduce the variability in the microstructure and mechanical properties. The laser power at each point of a certain layer needs to be controlled (increased/decreased by certain amount compared to previous layer) to achieve this homogeneity in melt pool temperature. The fundamental challenge of implementing a model predictive control system for such an application is the computational resources required for forward prediction (involving finite element or computational fluid dynamics models) and inverse prediction, as illustrated in Figure 3. The INN trainer aims to provide a reliable and memory-efficient surrogate model that can replace mechanistic models in predictive control systems. In this application, the database is generated by fusing experiments and computational methods (see SI Section 2.6 for details) for aluminum alloy. The online control loop uses the trained INN models to inform the manufacturing machine. There are two models: forward model and inverse model.
@@ 3.2 Calibration of Heat Source Parameters for Additive Manufacturing
where $z_{top}$ is the z-coordinate of the current surface of the powder bed and $\mathbf{1}$ is the indicator function. The notation follows Eq. 14, and $r_{b}$ is the Gaussian profile’s standard deviation that controls the beam’s width. Following reference [10], the heat source parameters $d,\eta,r_{b}$ are controlled by the calibration parameters $p_{1}$, $p_{2}$, $p_{3}$ such that $d=p_{1}\frac{P}{V}RHF^{2}$, $\eta=p_{2}\frac{P}{V}RHF^{2}$, and $r_{b}=p_{3}\frac{P}{V}RHF^{2}$. Here, $V$ is the laser scan speed and $RHF$ is a predetermined residual heat factor based on the laser toolpath [51].
This calibration problem’s major challenge is achieving a highly accurate surrogate model given a sparse set of training data. INNs outperform MLPs in solving these data-scarce problems because they are rooted in the interpolation theory. As a demonstration, an INN with 1760 trainable parameters is trained on 28 sets of meltpool temperature data generated by the finite difference method (FDM) [18] using 4 different power-scanning speed set-ups. Later, this trained INN is used as a surrogate to optimize the three parameters ($p_{1},p_{2},p_{3}$) given the experimental data. To show the superiority of INNs over traditional neural networks in such data-scarce applications, a 3-layer MLP with 2689 trainable parameters (roughly the same order of magnitude as the INN) is trained using the same data fed to the INN. In practice, we observed that INN converges well within 100 epoch, while MLP requires over 1000 epoch to converge to a test MSE 10 times larger than INN given all other setup the same. Details on this training and calibration process are provided in SI Section 3.1.
The INN- and MLP-based calibrations are repeated 50 times to provide a statistical comparison (see Figure 3). Even with only 28 sets of simulation data, the calibrated parameters using the INN produce a mean melt pool temperature within a 7.9% difference of the experimental data. In contrast, the parameters calibrated from the MLP surrogate produce a difference of 44. 1%, which is unsatisfactory. This is expected since MLP is a general model requiring a large amount of data to achieve a reliable representation of the problem. On the other hand, the INN achieves a much lower difference, showing that the model can learn underlying physics even with a scarce dataset.
3.3 Solving High Dimensional Space-Time-Parameter Heat Equation
@@
The surrogate model $f$ can also be obtained using the classical data-driven approach. In the offline stage, training data can be generated by repeatedly running numerical solvers (such as finite difference method (FDM) and variational multiscale method (VMS)) in the Space-Time domain with different sets of parameters $\mathbf{{p}}$. Then a surrogate model is trained with the data. The standard data-driven approach suffers from the curse of dimensionality when the parametric space is high-dimensional. As a result, the data-driven approaches face three challenges at the offline stage: expensive computation for repetitive data generation, huge memory cost for running full-scale simulations, and excessive disk storage of offline training data. We estimate the time and storage required for the data-driven approach with FDM and VMS, as shown in Table 2 and illustrated in Figure 2.
Unlike the data-driven approaches, INN solvers treat the parameters $\mathbf{p}$ as additional parametric inputs. As a result, INNs can directly obtain a parametric surrogate model $f=u(\mathbf{x},t,\mathbf{p})$ directly from the governing equation without going through the cumbersome offline data generation and surrogate model training process. This is particularly beneficial to high-dimensional parametric surrogate models. Moreover, our results prove the INN solution (or the S-T-P surrogate model) achieves an exceptionally high accuracy of $R^{2}$ = 0.9969, considering that most data-driven approaches for S-T-P problems trained on numerical simulation data suffer from low model accuracy, typically below $R^{2}=0.9$ [52]. It is worth mentioning that the physics-informed neural networks (PINNs) can handle the S-T-P problem similar to the INN solver. However, the number of collocation points scales exponentially with the number of inputs, and our preliminary study revealed that PINN fails to converge for the same 8-dimensional S-T-P problem after consuming considerable computational resources (see SI Section 4.4). The performance comparison of INN versus data-driven methods is shown in Table. 2.
@@
4 Conclusions
This article demonstrates that the interpolating neural network (INN) can train, solve, and calibrate scientific and engineering problems that are extremely challenging or prohibitive for existing numerical methods and machine learning models. The keys to INN’s success are 1) it interpolates the graph nodes’ values with well-established interpolation theories (not interpolating training data with a black-box network that often causes overfitting issues) and 2) it leverages the tensor decomposition (TD) to resolve curse of dimensionality. Due to the significantly reduced trainable parameters without losing approximation capabilities, INN can train/solve/calibrate faster with less memory requirement, and achieve extremely high model accuracy and fast inference time.
As this is the original article on INNs, we have myriads of research questions that may gather huge research interests across various areas including but not limited to scientific machine learning, applied mathematics, computer vision, and data science. To mention a few:
  • Depending on the problem, INN’s superior training efficiency may cause overfitting issues. The ensemble training (i.e., training on random subsets of the training data) might be a way to enhance the generality of the model for image problems or noisy data.
  • Multi-resolution aspect (i.e., a varying number of segments across TD modes) of INNs needs to be investigated. See SI 1.4 for further discussions.
  • Similar to the INN solvers, the convergence of INN trainers can be studied mathematically.
@@ 4 Conclusions
  • The superior performance of the INN solver may facilitate multiscale analysis in a vast parametric space. One can integrate the S-T-P INN solver with a concurrent multiscale analysis framework such as self-consistent clustering analysis [53, 54] and open a new direction towards parameterized multiscale analysis.
@@
  • Although Tucker decomposition has been introduced as a special case 1 in Section 2, the numerical experiments conducted in this article mostly focus on the CP decomposition. Theoretically, Tucker decomposition has a larger approximation space than CP decomposition. Further studies on the numerical aspects of the Tucker decomposition are needed.
@@ Acknowledgments
We express our sincere gratitude to Dr. Gino Domel, Mr. Joseph P. Leonor, Mr. Stefan Knapik, and Mr. Vispi Karkaria from Northwestern University for their invaluable comments and support throughout this work. We also acknowledge the following grants: DEVCOM Army Research Laboratory under Cooperative Agreement Numbers W911NF-20-2-0292 and W911NF-21-2-02199 to J. Cao and the startup funds provided by Virginia Polytechnic Institute and State University to S. Saha. C. Park acknowledges the internship opportunity with the Freudenberg Group, within the team managed by Dr. Vishwanath Hegadekatte.

### v5

@@ 3 Discussion
Now we illustrate how Engineering Software 2.0 enabled with INNs can be applied to advance various facets of metal additive manufacturing (AM). Note that the INN with CP decomposition will be referred to as INN in this section for brevity. Laser powder bed fusion (L-PBF) is a type of metal AM in which metal powders are layered and melted by a laser beam to create objects that match the provided CAD designs [41]. The added flexibility of L-PBF comes with a huge design space and spatially varying microstructure after manufacturing. Therefore, considerable research is devoted to employing in situ monitoring data for real-time control of L-PBF processes to minimize the variability and optimize the target properties [42]. Computational modeling of L-PBF spans the simulation of the manufacturing process with a high dimensional design space (solving), the identification of model parameters from sparse experimental data (calibrating), and the online monitoring and control of the L-PBF (training).
@@ 3.1 Real-time online control of laser powder bed fusion additive manufacturing
The first application is fully data-driven, where the data are generated from numerical methods and the INN trainer is used to develop an online control system for L-PBF AM. The goal is to maintain a homogeneous temperature of the melt pool (i.e., a molten region near the laser spot) across each layer using real-time sensor data, under the hypothesis that a homogeneous melt pool temperature across each layer of material deposition will reduce the variability in the microstructure and mechanical properties. The laser power at each point of a certain layer needs to be controlled (increased/decreased by a certain amount compared to the previous layer) to achieve this homogeneity in melt pool temperature. The fundamental challenge of implementing a model predictive control system for such an application is the computational resources required for forward prediction (involving finite element or computational fluid dynamics models) and inverse prediction, as illustrated in Figure 4, the row for training. The INN trainer aims to provide a reliable and memory-efficient surrogate model that can replace mechanistic models in predictive control systems. In this application, the database is generated by fusing experiments and computational methods (see SI Section 2.6 for details) for aluminum alloy. The online control loop uses the trained INN models to inform the manufacturing machine. There are two models: forward model and inverse model.
@@ 3.2 Calibration of heat source parameters for additive manufacturing
where $z_{top}$ is the z-coordinate of the current surface of the powder bed and $\mathbf{1}$ is the indicator function that returns 1 if the condition is met and 0 otherwise. The notation follows Eq. 4, and $r_{b}$ is the Gaussian profile’s standard deviation that controls the beam’s width. Following reference [10], the heat source parameters $d,\eta,r_{b}$ are controlled by the calibration parameters $p_{1}$, $p_{2}$, $p_{3}$ such that $d=p_{1}\frac{P}{V}RHF^{2}$, $\eta=p_{2}\frac{P}{V}RHF^{2}$, and $r_{b}=p_{3}\frac{P}{V}RHF^{2}$. Here, $V$ is the laser scan speed and $RHF$ is a predetermined residual heat factor based on the laser toolpath [45].
The major challenge of this problem is to achieve a highly accurate surrogate model given a sparse set of training data. INNs outperform MLPs in training these scarce data because they are rooted in the interpolation theory. As a demonstration, an INN with 1760 trainable parameters is trained on 28 sets of meltpool temperature data generated by the finite difference method (FDM) [24] using 4 different power-scanning speed set-ups. Later, this trained INN is used as a surrogate to optimize the three parameters ($p_{1},p_{2},p_{3}$) given the experimental data. For comparison, a 3-layer MLP with 2689 trainable parameters (roughly the same order of magnitude as the INN) is trained using the same data. We observed that INNs converge well within 100 epochs, while MLPs require over 1000 epochs to converge to a test MSE 10 times larger than the INN given all other setups the same. Details on this training and calibration process are provided in SI Section 3.1.
The INN- (definition of the INN calibrator can be found in Section 5.5 and MLP-based calibrations are repeated 50 times to provide a statistical comparison (see Figure 4). Even with only 28 sets of simulation data, the calibrated parameters using the INN produce a mean melt pool temperature within a 7.9% difference of the experimental data. In contrast, the parameters calibrated from the MLP surrogate produce a difference of 44. 1%, which is unsatisfactory. This is expected since MLP is a general model requiring a large amount of data to achieve a reliable representation of the problem. On the other hand, the INN achieves a much lower difference, showing that the model can learn underlying physics even with a scarce dataset.
3.3 Solving high dimensional Space-Time-Parameter (S-T-P) heat equation
@@
The surrogate model $f$ can also be obtained using the standard data-driven approach. In the offline stage, training data are generated by repeatedly running numerical solvers (such as FDM or variational multiscale method (VMS)) in the Space-Time domain with different sets of parameters $\mathbf{{p}}$. Then a surrogate model is trained with the data. The data-driven approach suffers from the curse of dimensionality when the parametric space is high-dimensional, resulting in expensive computation for repetitive data generation, huge memory costs for running full-scale simulations, and excessive disk storage of offline training data. We estimate the time and storage required for the data-driven approaches with FDM and VMS, as shown in Table 1 and illustrated in Figure 3(a).
Unlike the data-driven approaches, INN solvers treat the parameters $\mathbf{p}$ as additional parametric inputs. As a result, INNs obtain a parametric surrogate model $f=u(\mathbf{x},t,\mathbf{p})$ directly from the governing equation without going through the cumbersome offline data generation and surrogate model training. Moreover, our results prove the INN solution (or the S-T-P surrogate model) achieves an exceptionally high accuracy of $R^{2}$ = 0.9969, considering that most data-driven approaches for S-T-P problems trained on numerical simulation data suffer from low model accuracy, typically below $R^{2}=0.9$ [46]. It is worth mentioning that the physics-informed neural networks (PINNs) can handle the S-T-P problem similar to the INN solver. However, the number of collocation points scales exponentially with the number of inputs, and our preliminary study revealed that PINN fails to converge for the same 8-dimensional S-T-P problem after consuming considerable computational resources (see SI Section 4.4). The performance comparison of INN versus data-driven methods is shown in Table. 1.
@@
4 Conclusion
This article demonstrates that interpolating neural networks (INNs) can train, solve, and calibrate scientific and engineering problems that are extremely challenging or prohibitive for existing numerical methods and machine learning models. The keys to INN’s success are 1) it interpolates the graph nodes’ values with well-established interpolation theories and 2) it leverages the tensor decomposition (TD) to resolve the curse of dimensionality. Due to the reduced number of parameters without compromising predictive accuracy, INNs become an efficient substitute for MLPs or PINNs.
There are numerous research questions that could generate significant interest across various fields, including, but not limited to, scientific machine learning, applied mathematics, computer vision, and data science. To mention a few:
  • Depending on the problem, INN’s superior training efficiency may overfit data. The ensemble training (i.e., training on random subsets of the training data) might be a way to enhance the generality of the INN model for real-world data.
  • Multi-resolution aspect (i.e., a varying mesh resolution across TD modes) of INNs needs to be investigated. See SI 1.4 for further discussions.
  • The interpretability of INNs should be investigated. INNs may address the challenge of model-based interpretability (a terminology defined from [47]), which involves developing models that are both simple enough to be easily understood by developers and capable of maintaining high predictive accuracy.
  • Similar to INN solvers, the convergence of INN trainers can be studied both mathematically and numerically.
@@ 4 Conclusion
  • The superior performance of INN solvers may facilitate multiscale analysis in a vast parametric space. One can integrate the S-T-P INN solver with a concurrent multiscale analysis framework such as self-consistent clustering analysis [48, 49] and open a new direction towards parameterized multiscale analysis.
@@
  • Although Tucker decomposition is introduced in Section 5.1, the numerical experiments conducted in this article mostly focus on the CP decomposition. Theoretically, Tucker decomposition has a larger approximation space than CP decomposition. Further studies on the numerical aspects of the Tucker decomposition are needed.
@@
5 Methods
5.1 Special case 1: regular mesh with Tucker decomposition
One of the widely used TD methods is Tucker decomposition [33, 34]. It approximates a high-order tensor with $J_{i}$ nodes in $i$-th dimension as a tensor contraction between dimension-wise matrices and a core tensor $\bm{\mathcal{G}}$, which has the same order of the original tensor but with smaller nodes $M_{i}$ ($<J_{i}$). The $M_{i}$ is often called a "mode" to distinguish it from the original node $J_{i}$.
Consider a three-input ($I=3$) and one output ($L=1$) system, and assume the input domain is discretized with $J=J_{1}\times J_{2}\times J_{3}$ nodes, as shown in the special case box of Figure 1. To facilitate tensor notation, the nodal values will be denoted with left/right super/sub scripts, $\prescript{m}{i}{u}^{(j)}_{l}$, where $i\in\mathbb{N}^{I}$ is the input index, $m\in\mathbb{N}^{M_{i}}$ is the mode index, $l\in\mathbb{N}^{L}$ is the output index, and $j\in\mathbb{N}^{J_{i}}$ is the nodal index.
where the core tensor $\bm{\mathcal{G}}\in\mathbb{R}^{M_{1}\times M_{2}\times M_{3}}$ is a trainable full matrix typically smaller than the original tensor that compresses the data. Here, $\bm{\mathcal{A}}\times^{b}_{a}\bm{\mathcal{B}}$ denotes the tensor contraction operation between the $a$-th dimension of tensor $\bm{\mathcal{A}}$ and the $b$-th dimension of tensor $\bm{\mathcal{B}}$. The $\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})\in\mathbb{R}^{M_{i}\times 1}$ is one-dimensional (1D) interpolated output of $i$-th input dimension over $M_{i}$ modes represented as:
$\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})=\begin{bmatrix}\prescript{1}{i}{u}^{(1)}&\prescript{1}{i}{u}^{(2)}&\cdots&\prescript{1}{i}{u}^{(J_{i})}\\ \prescript{2}{i}{u}^{(1)}&\prescript{2}{i}{u}^{(2)}&\cdots&\prescript{2}{i}{u}^{(J_{i})}\\ \vdots&\vdots&\vdots&\vdots\\ \prescript{M_{i}}{i}{u}^{(1)}&\prescript{M_{i}}{i}{u}^{(2)}&\cdots&\prescript{M_{i}}{i}{u}^{(J_{i})}\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(x_{i})\\ N^{(2)}(x_{i})\\ \vdots\\ N^{(J_{i})}(x_{i})\end{bmatrix}=\prescript{}{i}{\bm{U}}\times^{1}_{2}\prescript{}{i}{\mathcal{X}(x_{i})}=\prescript{}{i}{\bm{U}}\prescript{}{i}{\mathcal{X}(x_{i})},$ (7)
where $\prescript{}{i}{\bm{U}}\in\mathbb{R}^{M_{i}\times J_{i}}$ and $\prescript{}{i}{\bm{\mathcal{X}(x)}}\in\mathbb{R}^{J_{i}}$. As illustrated in the special case box of Figure 1, the message passing (blue arrow) only happens in the axial directions, yielding 1D interpolation functions: $N^{(J_{i})}(x_{i})$.
When there are more than one output $(L>1)$, the interpolated field becomes a vector of $L$ elements:
5.2 Special case 2: regular mesh with CP decomposition
where $\mathcal{J}\prescript{m}{i}{\bm{u}}(x_{i})\in\mathbb{R}^{L}$ and $\odot$ represent multiplication in elements. In CP decomposition, the core tensor is no longer trainable, thus the total trainable parameter becomes $ML\sum^{I}_{i}{J_{i}}$.
It is important to note that the interpolation function $N^{(j)}(x_{i})$ in Eq.7 is one-dimensional. In other words, both Tucker and CP decomposition replace a high-dimensional interpolation with one-dimensional interpolations that dramatically reduce the number of trainable parameters (or DoFs). As given in Table 2, the number of trainable parameters of the full interpolation scales exponentially with the discretization $J_{i}$, whereas the Tucker and CP decomposition scale linearly with $J_{i}$. Considering the trainable parameters of multi-layer perceptron (MLP) scales quadratically with the number of hidden neurons, INNs with tensor decomposition (TD) may dramatically reduce the model complexity and computing requirements.
However, the reconstructed basis using TD loses expressibility because of the separated variables. There is no cross-term in the functional space of TD. Nevertheless, we can transform this into an advantage by enriching the 1D approximations with adaptive activation (see SI Section 2.2) for the Q-step message passing. Therefore, we can find a hybrid architecture that preserves some of the advantages of numerical methods (i.e., FEM) and neural network methods while acknowledging that it loses something from both.
Full interpolation Tucker decomposition CP decomposition
5.3 Loss function of INN trainer
A general INN forward propagation is provided in Eq.3 as a tensor contraction. INN can be used for data training as any other neural network architecture. While MLPs optimize weights and biases during training, INN trainers optimize nodal values $\bm{U}=\{\bm{u}^{(j)}\}_{j=1,\cdots,J}$ (and, if needed, nodal coordinates ($\bm{X}=\{\bm{x}^{(j)}\}_{j=1,\cdots,J}$)) under a given loss function and training data. Consider a regression problem with $K$ labeled data: $(\bm{x}^{*}_{k},\bm{u}^{*}_{k})$, $k=1,\cdots,K$; and $\bm{x}^{*}_{k}\in\mathbb{R}^{I},\bm{u}^{*}_{k}\in\mathbb{R}^{L}$. The superscript * denotes the data. A mean squared error (MSE) loss function for this regression problem is defined as:
Finally, an optimization is formulated as follows.
$\underset{\bm{U},\bm{X}}{\text{minimize}}\ loss(\bm{U},\bm{X}).$ (11)
5.4 Loss function of INN solver
An INN solver generalizes classical numerical methods such as FEM and meshfree, as well as model order reduction methods such as proper generalized decomposition (PGD) [50] and tensor decomposition (TD) [20, 51]. Here, we introduce a formulation with CP decomposition to solve the generalized space $(\bm{x})$ time $(t)$ parameter $(\bm{\theta})$ (S-T-P) PDE, whose computational cost is prohibitively high for most numerical methods and machine learning approaches [50].
Consider a parameterized space-time PDE:
$\mathcal{L}\bm{u}(\bm{x},t,\bm{\theta})=0,$ (12)
[…]
where $\delta\bm{u}$ is the weight function and can be defined using Galerkin, Petrov-Galerkin, collocation or quadratic formulation. Due to the Kronecker delta property of INN interpolation functions, Dirichlet boundary conditions and initial conditions can be strongly imposed. As a result, only the weighted sum residual is considered in Eq. 13 as a loss function. This distinguishes INNs from most data-driven PDE solvers that weakly impose these conditions [55]. See Section 4 of the SI for detailed derivations of the loss function and solution scheme. INN solvers applied to metal additive manufacturing and the linear elastic solid mechanics problem are introduced in Section 3.3 and SI Section 4, respectively.
5.5 Loss function of INN calibrator
[…]
5.6 1D Poisson’s equation with error estimator
We borrow the manufactured problem first introduced in [14]. The 1D Poisson’s equation is defined as:
$\Delta u(x)+b(x)=0,\quad x\in[0,10],$ (16)
$u(x=0)=u(x=10)=0,$
where $\Delta$ is the Laplace operator. The manufactured solution becomes
$u(x)=(e^{-\pi(x-2.5)^{2}}-e^{-6.25\pi})+2(e^{-\pi(x-7.5)^{2}}-e^{-56.25\pi})-\frac{e^{-6.25\pi}-e^{-56.25\pi}}{10}x$ (17)
when the body force is given as:
$b(x)=-\frac{4\pi^{2}(x-2.5)^{2}-2\pi}{e^{\pi(x-2.5)^{2}}}-\frac{8\pi^{2}(x-7.5)^{2}-4\pi}{e^{\pi(x-7.5)^{2}}}.$ (18)
The $H^{1}$ norm error estimator is defined as:
[…]
where $u^{h}(x)$ is the approximated solution.
@@ Acknowledgments
We express our sincere gratitude to Dr. Gino Domel, Mr. Joseph P. Leonor, Mr. Stefan Knapik, and Mr. Vispi Karkaria from Northwestern University for their invaluable comments and support throughout this work.
Code availability
The computer codes used for the benchmarking can be found at https://github.com/hachanook/pyinn. Codes for additive manufacturing examples can be provided upon a reasonable request.